Calculate the pH of 0.010 M Ca(CH3CO2)2 (calcium acetate) solution.

A. 2.00

B. 5.47

C. 5.62

D. 8.37

E. 8.52
Answer #1

[CH3CO2-] = 2*[Ca(CH3CO2)2]

= 2*0.010 M

= 0.020 M

Ka of CH3CO2H is 1.8*10^-5

lets find Kb for CH3CO2-

use:

Kb = Kw/Ka

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

Kb = (1.0*10^-14)/Ka

Kb = (1.0*10^-14)/1.8*10^-5

Kb = 5.556*10^-10

CH3CO2- dissociates as

CH3CO2- + H2O -----> CH3CO2H + OH-

0.02 0 0

0.02-x x x

Kb = [CH3CO2H][OH-]/[CH3CO2-]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((5.556*10^-10)*2*10^-2) = 3.333*10^-6

since c is much greater than x, our assumption is correct

so, x = 3.333*10^-6 M

use:

pOH = -log [OH-]

= -log (3.333*10^-6)

= 5.48

use:

PH = 14 - pOH

= 14 - 5.48

= 8.52

Answer: 8.52